# A non-degeneracy theorem for interacting electrons in one dimension

### v1

@@ 5.2. Unitary reduction to the simplex
and therefore $T$ is an isometry. Moreover, a similar calculation shows that the restriction map $\Psi\mapsto T^{-1}\Psi=\sqrt{N!}\Psi\rvert_{S_{N}}$ is an isometry from $\mathrm{H}^{1}(I_{N})\cap\mathcal{H}_{N}$ to $\mathrm{H}^{1}(S_{N};\Gamma_{\rm int})$. As $T^{-1}$ maps $\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}$ to $\mathrm{H}^{1}_{L}(S_{N})$ by definition, it follows that $T:\mathrm{H}^{1}_{L}(S_{N})\rightarrow\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}$ is surjective, which completes the proof. ∎
As the operator $T$ in (5.6) is an isometry from $\mathrm{H}^{1}_{L}(S_{N})$ to $\mathrm{H}^{1}_{L}(I_{N})$ and from $\mathrm{L}^{2}(S_{N})$ to $\mathrm{L}^{2}(I_{N})$, by Lemma A.8, the form
$\widetilde{a}_{v,w}(\Psi,\Phi)$ $\coloneqq(T^{\#}a_{v,w})(\Psi,\Phi)=a_{v,w}(T\Psi,T\Phi)=\int_{I_{N}}\overline{\nabla(T\Psi)(x)}\bm{\cdot}\nabla(T\Phi)(x)\mathrm{d}x+v(\rho_{T\Psi,T\Phi})+w(\rho_{2,T\Phi,T\Phi}),$
$=\int_{S_{N}}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\Phi(x)\mathrm{d}x+v(\rho_{T\Psi,T\Phi})+w(\rho_{2,T\Phi,T\Phi}),\quad\mbox{for $\Psi,\Phi\in\mathrm{H}^{1}_{L}(S_{N})$,}$ (5.8)
@@
$\widetilde{H}^{L}_{N}(v,w)\coloneqq T^{\#}H_{N}^{L}(v,s)$ (5.9)
as the unique self-adjoint operator associated to the form $\widetilde{a}_{v,w}$ with form domain $\mathrm{H}^{1}_{L}(S_{N})$. Moreover, since $T$ is an isometric isomorphism between $\mathrm{H}^{1}_{L}(S_{N})$ and $\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}$, the operator $H_{N}^{L}(v,w)$ is unitarily equivalent to $\widetilde{H}_{N}^{L}(v,w)$. In particular, we can now apply Theorem 5.3 to obtain the following result.
@@
Let $v\in\mathcal{V}$, $w\in\mathcal{W}$, and $L\subset\mathrm{H}^{1/2}(\partial I_{N})$ be a closed subspace such that
$\mathrm{H}^{1}_{L}(S_{N})=\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}\}$
is reality and positivity preserving. Then the self-adjoint operator $H_{N}^{L}(v,w)$ defined in Lemma 4.1 has a non-degenerate ground-state and the unique (up to a global phase) ground-state wave-function $\Psi$ satisfies $\Psi(x)>0$ for a.e. $x\in S_{N}$.
@@
As $T$ is an isometry, the ground-state of the reduced operator $\widetilde{H}^{L}_{N}(v,w)$ defined in (5.9) is mapped to the ground-state of $H^{L}_{N}(v,w)$ via the map $\Psi\mapsto T\Psi$. Hence, it suffices to show that $H_{N}^{L}(v,w)$ has a non-degenerate and strictly positive ground-state. In particular, it suffices to show that $\widetilde{H}^{L}_{N}(v,w)$ satisfy the assumptions from Theorem 5.3.
@@
$\lim_{n\rightarrow\infty}\sup_{\Psi,\Phi\in\mathrm{H}^{1}(I^{N})\setminus\{0\}}\frac{|v_{n}(\rho_{\Psi,\Phi})+w_{n}(\rho_{2,\Psi,\Phi})-v(\rho_{\Psi,\Phi})-w(\rho_{2,\Psi,\Phi})|}{\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}}\leq\lim_{n\rightarrow\infty}\lVert v_{n}-v\rVert_{-1,2}+\lVert w_{n}-w\rVert_{-1,p}=0.$
As $T:\mathrm{H}^{1}_{L}(S_{N})\rightarrow\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}$ is continuous, the forms $T^{\#}v_{n}$ and $T^{\#}w_{n}$, defined as
$(T^{\#}v_{n})(\Psi,\Phi)=v_{n}(\rho_{T\Psi,T\Phi})\quad\mbox{and}\quad(T^{\#}w_{n})(\Psi,\Phi)=w_{n}(\rho_{2,T\Psi,T\Phi}),$
@@
$\lim_{n\rightarrow\infty}\sup_{\Psi,\Phi\in\mathrm{H}^{1}_{L}(S_{N})\setminus\{0\}}\frac{|(T^{\#}v_{n}+T^{\#}w_{n}-T^{\#}v-T^{\#}w)(\Psi,\Phi)|}{\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}}=0.$ (5.10)
@@
$(T^{\#}V)(\Psi,\Phi)$ $=V(T\Psi,T\Phi)=\int_{I_{N}}V(x)\overline{T\Psi(x)}T\Phi(x)\mathrm{d}x=\frac{1}{N!}\sum_{\sigma,\tau\in\mathcal{P}_{N}}\int_{\sigma(S_{N})\cap\tau(S_{N})}V(x)\overline{\Psi(\sigma^{-1}x)}\Phi(\tau^{-1}x)\mathrm{d}x,$
$=\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}\int_{S_{N}}V(\sigma x)\overline{\Psi(x)}\Phi(x)\mathrm{d}x=\int_{S_{N}}\widetilde{V}(x)\overline{\Psi(x)}\Phi(x)\mathrm{d}x.$
In particular, this implies that the pull-back of any form on $I_{N}$ generated by a bounded multiplicative potential is a form in $S_{N}$ generated by a bounded multiplicative potential. As the forms $v_{n}$ and $w_{n}$ are generated by the multiplicative potentials
@@
the forms $T^{\#}(v_{n}+w_{n})$ are also generated by multiplicative potentials. Thus, by estimate (5.10), we have found a sequence of multiplicative potentials approximating $T^{\#}v+T^{\#}w$. We can therefore apply Theorem 5.3 to complete the proof. ∎
@@
$\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}\}\subset\mathrm{H}^{1}(S_{N})$ (5.11)
@@
$\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}\}\subset\mathrm{H}^{1}(S_{N})$ (5.12)
is also reality and positivity preserving.
@@
That both spaces are reality preserving is clear since $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)$ with $\alpha\in\mathbb{R}$ and $\mathcal{H}_{N}$ are both reality preserving. For the positivity preserving part, we first deal with the space in (5.11).
In this case, the proof follows from two observations. First, by antisymmetry, the boundary trace of any function $\Psi$ in $\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ must actually vanish on the (possibly larger) symmetrized set $\Gamma_{N}=\cup_{\sigma\in\mathcal{P}_{N}}\sigma(\Gamma)$. In particular, the restriction $\Psi\rvert_{S_{N}}$ vanishes on the intersection $\Gamma_{N}\cap\partial S_{N}$. The second observation is that $(\gamma\Psi)_{+}=\gamma(\Psi_{+})$ because this holds for continuous functions and taking the positive part is (e.g., by [MM79, Theorem 1]) a continuous operation in $\mathrm{H}^{1}$. Therefore, $(\Psi\rvert_{S_{N}})_{+}\in\mathrm{H}^{1}(S_{N};\Gamma_{N}\cap\partial I_{N})$. In particular, the boundary trace of the extension $T(\Psi_{+}\rvert_{S_{N}})$ vanishes on the symmetrized set $\Gamma_{N}$ and therefore on $\Gamma$. Thus $T(\Psi_{+}\rvert_{S_{N}})$ belongs to $\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ and satisfies $T(\Psi_{+}\rvert_{S_{N}})\rvert_{S_{N}}=\frac{1}{\sqrt{N!}}\Psi_{+}\rvert_{S_{N}}$. This shows that $\Psi_{+}$ belongs to the space in (5.11) and therefore this space is positivity preserving.
@@
$Y\coloneqq\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}\}$
is positivity preserving. Indeed, if $Y$ is positivity preserving, then the result follows because the intersection of two positivity preserving spaces is positivity preserving, and
$\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}\}=Y\cap\mathrm{H}^{1}(S_{N};\Gamma_{N}\cap\partial S_{N}).$
To prove that $Y$ is positivity preserving, we first claim that
$Y=\{\Psi\in\mathrm{H}^{1}(S_{N};\Gamma_{\rm int}):\Psi(0,x^{\prime})-\alpha(-1)^{N-1}\Psi(x^{\prime},1)=0\quad\mbox{for a.e. $x^{\prime}\in S_{N-1}$}\}.$ (5.13)
@@
and therefore $Y$ is positivity preserving.
To prove the claim, we first note that, by the definition of $\mathrm{H}^{1}_{\alpha}(I_{N})$ and antisymmetry, any $\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}$ satisfies
@@
for any $x^{\prime}=(x_{1},...,x_{N-1})\in I_{N-1}$; therefore the inclusion $\subset$ in (5.13) holds. To prove the opposite inclusion, we observe that for any $\Psi\in\mathrm{H}^{1}(S_{N};\Gamma_{\rm int})$ satisfying $\Psi(0,x^{\prime})=(-1)^{N-1}\alpha\Psi(x^{\prime},1)$ for a.e. $x^{\prime}\in S_{N-1}$, we have
$(T\Psi)(0,\sigma x^{\prime})$ $=\operatorname{sign}(\sigma)\Psi(0,x^{\prime})=\operatorname{sign}(\sigma)(-1)^{N-1}\alpha\Psi(x^{\prime},1)=(-1)^{N-1}\alpha(T\Psi)(\sigma x^{\prime},1)$
@@
for any $\sigma\in\mathcal{P}_{N-1}$ and a.e. $x^{\prime}\in S_{N-1}$. As $\{\sigma(S_{N-1})\}_{\sigma\in\mathcal{P}_{N-1}}$ covers $I_{N-1}$ up to finitely many hyperplanes of dimension $N-2$, equation (5.14) holds for almost every $x^{\prime}\in I_{N-1}$. Thus $T\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N})$ and satisfies $(T\Psi)\rvert_{S_{N}}=\frac{1}{\sqrt{N!}}\Psi$, which proves the inclusion $\supset$ in (5.13) and completes the proof. ∎
@@
Let $\alpha\in\mathbb{R}\setminus\{0\}$ and $v\in\mathcal{V}$, then by Lemma 4.4, the eigenfunctions of the operator $H_{N}^{\alpha}(v,0)$ are given by linear combination of Slater determinants of the eigenfunctions of $h_{\alpha}(v)=-\Delta_{\alpha}+v$. In particular, if we denote by $\{\psi_{j}\}_{j\in\mathbb{N}}$ and $\{\lambda_{j}\}_{j\in\mathbb{N}}$ the eigenfunctions and eigenvalues of $h_{\alpha}(v)$ ordered in non-decreasing order, then $H_{N}(v,0)$ has a non-degenerate ground-state if and only if $\lambda_{N}<\lambda_{N+1}$. Thus by Theorem 2.3, we have $\lambda_{N}<\lambda_{N+1}$ for any $N\in\mathbb{N}$ satisfying (2.4), in particular, for $N=2k-1$ if $\alpha\geq 0$ and $N=2k$ for $\alpha\leq 0$. This proves (2.7). The cases of local (or separable) boundary conditions in (2.8) follows from the same argument.

### v2

@@ 4.2. Unitary reduction to the simplex
and therefore $T$ is an isometry. Moreover, a similar calculation shows that the restriction map $\Psi\mapsto T^{-1}\Psi=\sqrt{N!}\Psi\rvert_{S_{N}}$ is an isometry from $\mathrm{H}^{1}(I_{N})\cap\mathcal{H}_{N}$ to $\mathrm{H}^{1}(S_{N};\Gamma_{\rm int})$. As $T^{-1}$ maps $\mathcal{Q}_{N}(B)=\mathrm{H}^{1}_{B}(I_{N})\cap\mathcal{H}_{N}$ to $\mathcal{Q}_{N}(B;S_{N})$ by definition, it follows that $T:\mathcal{Q}_{N}(B;S_{N})\rightarrow\mathcal{Q}_{N}(B)$ is surjective, which completes the proof. ∎
As the operator $T$ in (4.6) is an isometry from $\mathcal{Q}_{N}(B;S_{N})$ to $\mathcal{Q}_{N}(B)$, the form
$\widetilde{\mathfrak{a}}_{v,w}(\Psi,\Phi)$ $\coloneqq(T^{\#}\mathfrak{a}_{v,w})(\Psi,\Phi)=\mathfrak{a}_{v,w}(T\Psi,T\Phi)=\int_{I_{N}}\overline{\nabla(T\Psi)(x)}\boldsymbol{\cdot}\nabla(T\Phi)(x)\mathrm{d}x+v(\rho_{T\Psi,T\Phi})+w(\rho^{(2)}_{T\Phi,T\Phi})$,
$=\int_{S_{N}}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\Phi(x)\mathrm{d}x+v(\rho_{T\Psi,T\Phi})+w(\rho^{(2)}_{T\Phi,T\Phi}),\quad\mbox{for $\Psi,\Phi\in\mathcal{Q}_{N}(B;S_{N})$,}$ (4.8)
@@
$\widetilde{H}^{B}_{N}(v,w)\coloneqq T^{\#}\left(H_{N}^{B}(v,w)\right)=T^{-1}H_{N}^{B}(v,w)T$ (4.9)
as the unique self-adjoint operator associated to the form $\widetilde{\mathfrak{a}}_{v,w}$ with form domain $\mathcal{Q}_{N}(B;S_{N})$. Moreover, this operator is unitarily equivalent to $H^{B}_{N}(v,w)$ and we can apply Theorem 4.6 to obtain the following result.
@@
Let $v\in\mathcal{V}$, $w\in\mathcal{W}$, and $B\subset\mathrm{H}^{1/2}(\partial I_{N})$ be a closed subspace such that the space defined in (4.5),
$\mathcal{Q}_{N}(B;S_{N})=\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{B}(I_{N})\cap\mathcal{H}_{N}\},$
is reality and positivity preserving. Then the self-adjoint operator $H_{N}^{B}(v,w)$ defined via (2.1) (see also Lemma A.1) has a non-degenerate ground-state and the unique (up to a global phase) ground-state wave-function $\Psi$ satisfies $\Psi(x)>0$ for a.e. $x\in S_{N}$.
@@
By construction, the ground-state of $H^{B}_{N}(v,w)$ is given by $T\Psi$ where $\Psi$ is the ground-state of $\widetilde{H}^{B}_{N}(v,w)$. Hence, it suffices to show that $\widetilde{H}_{N}^{B}(v,w)$ has a non-degenerate and strictly positive ground-state. For this, note that $C^{\infty}_{c}(S_{N})\subset\mathcal{Q}_{N}(B;S_{N})$. In particular, by Proposition 3.9, the form domain $\mathcal{Q}_{N}(B;S_{N})$ is of the form $\mathrm{H}^{1}_{B^{\prime}}(S_{N})$ for a suitable closed subspace $B^{\prime}\subset\mathrm{H}^{\frac{1}{2}}(\partial S_{N})$. Hence, it suffices to show that $\widetilde{H}^{B}_{N}(v,w)$ satisfy the assumptions from Theorem 4.6.
@@
$\lim_{n\rightarrow\infty}\sup_{\Psi,\Phi\in\mathrm{H}^{1}(I^{N})\setminus\{0\}}\frac{|v_{n}(\rho_{\Psi,\Phi})+w_{n}(\rho^{(2)}_{\Psi,\Phi})-v(\rho_{\Psi,\Phi})-w(\rho^{(2)}_{\Psi,\Phi})|}{\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}}\leq\lim_{n\rightarrow\infty}\lVert v_{n}-v\rVert_{-1,2}+\lVert w_{n}-w\rVert_{-1,p}=0.$
As $T:\mathcal{Q}_{N}(B;S_{N})\rightarrow\mathcal{Q}_{N}(B)$ is continuous, the forms $T^{\#}v_{n}$ and $T^{\#}w_{n}$, defined as
$(T^{\#}v_{n})(\Psi,\Phi)=v_{n}(\rho_{T\Psi,T\Phi})\quad\mbox{and}\quad(T^{\#}w_{n})(\Psi,\Phi)=w_{n}(\rho^{(2)}_{T\Psi,T\Phi})$,
@@
$\lim_{n\rightarrow\infty}\sup_{\Psi,\Phi\in\mathcal{Q}_{N}(B;S_{N})\setminus\{0\}}\frac{|(T^{\#}v_{n}+T^{\#}w_{n}-T^{\#}v-T^{\#}w)(\Psi,\Phi)|}{\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}}=0.$ (4.10)
@@
$(T^{\#}V)(\Psi,\Phi)$ $=V(T\Psi,T\Phi)=\int_{I_{N}}V(x)\overline{(T\Psi)(x)}(T\Phi)(x)\mathrm{d}x$
$=\frac{1}{N!}\sum_{\sigma,\tau\in\mathcal{P}_{N}}\mathrm{sgn}(\sigma)\mathrm{sgn}(\tau)\int_{\sigma(S_{N})\cap\tau(S_{N})}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!V(x)\overline{\Psi(\sigma^{-1}x)}\Phi(\tau^{-1}x)\mathrm{d}x=\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}\int_{S_{N}}V(\sigma x)\overline{\Psi(x)}\Phi(x)\mathrm{d}x$
$=\int_{S_{N}}\widetilde{V}(x)\overline{\Psi(x)}\Phi(x)\mathrm{d}x,\quad\mbox{where}\quad\widetilde{V}(x_{1},...,x_{N})=\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}V(\sigma x).$
In particular, this implies that the pull-back of any form generated by a bounded potential in $I_{N}$ is a form generated by a bounded potential in $S_{N}$. As the forms $v_{n}$ and $w_{n}$ are generated by the bounded potentials
@@
the forms $T^{\#}(v_{n}+w_{n})$ are also generated by bounded potentials. Thus, by estimate (4.10), we have found a sequence of bounded multiplicative potentials approximating $T^{\#}v+T^{\#}w$. We can therefore apply Theorem 4.6 to complete the proof. ∎
Remark 4.9 (Bose-Fermi equivalence in 1D).
By considering the symmetric map
$T_{+}:\mathcal{Q}_{N}(B;S_{N})\rightarrow\mathrm{H}^{1}(I_{N})\cap\mathcal{H}^{\rm Bos}_{N},\qquad\qquad T_{+}(\Psi)\coloneqq\frac{1}{\sqrt{N!}}\sum_{\sigma\in\mathcal{P}_{N}}\sigma_{\#}(\Psi),$
where $\mathcal{H}^{\rm Bos}_{N}=\otimes^{N}_{\rm sym}\mathrm{L}^{2}(I)$ is the symmetric tensor-product, we obtain an isomorphism between $\mathcal{Q}_{N}(B;S_{N})$ and a subset of bosonic (symmetric) wavefunctions vanishing along the hyperplanes $\{x\in I_{N}:x_{j}=x_{k}\quad\mbox{for some $j\neq k$}\}$. This vanishing behaviour can be understood as the effect of a a limiting hard-core pairwise interaction potential $w(x,y)=w_{a}(x-y)=+\infty$ for $|x-y|\leq a$ and zero otherwise, in the case where $a\downarrow 0$. In particular, we note that the composition map
$T_{+}\circ T^{-1}(\Psi)(x_{1},...,x_{N})=\prod_{j>k}\mathrm{sign}(x_{j}-x_{k})\Psi(x_{1},...,x_{N})$
is precisely the isomorphism (sign map) between fermions and hard-core bosons explicitly constructed by Girardeau in [Gir60]. Therefore, we can recover the Fermi-Bose equivalence due to [Gir60] for a general class of interactions.
@@
$\mathcal{Q}_{N}(\Gamma;S_{N})\coloneqq\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}\}\subset\mathrm{H}^{1}(S_{N})$ (4.11)
@@
$\mathcal{Q}_{N}^{\alpha}(\Gamma;S_{N})\coloneqq\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}\}\subset\mathrm{H}^{1}(S_{N}),$ (4.12)
where $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)$ is defined in (2.6), is also reality and positivity preserving.
@@
That both spaces are reality preserving is clear since $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)$ with $\alpha\in\mathbb{R}$ and $\mathcal{H}_{N}$ are both reality preserving. For the positivity preserving part, we first deal with $\mathcal{Q}_{N}(\Gamma;S_{N})$.
In this case, the proof follows from two observations. First, by antisymmetry, the boundary trace of any function $\Psi$ in $\mathcal{Q}_{N}$ is anti-symmetric and must therefore vanish on the (possibly larger) symmetrized set $\Gamma_{N}=\cup_{\sigma\in\mathcal{P}_{N}}\sigma(\Gamma)$. In particular, the restriction $\Psi\rvert_{S_{N}}$ vanishes on the intersection $\Gamma_{N}\cap\partial S_{N}$. The second observation is that $(\gamma\Psi)_{+}=\gamma(\Psi_{+})$ because this holds for continuous functions and taking the positive part is (e.g., by [MM79, Theorem 1]) a continuous operation in $\mathrm{H}^{1}$. Therefore, $(\Psi\rvert_{S_{N}})_{+}\in\mathrm{H}^{1}(S_{N};\Gamma_{N}\cap\partial I_{N})$. In particular, the boundary trace of the extension $T(\Psi_{+}\rvert_{S_{N}})$ vanishes on the symmetrized set $\Gamma_{N}$ and therefore on $\Gamma$. Thus $T(\Psi_{+}\rvert_{S_{N}})$ belongs to $\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ and satisfies $T(\Psi_{+}\rvert_{S_{N}})\rvert_{S_{N}}=\frac{1}{\sqrt{N!}}\Psi_{+}\rvert_{S_{N}}$. This shows that $\Psi_{+}$ belongs to $\mathcal{Q}_{N}(\Gamma;S_{N})$, which implies that this space is positivity preserving.
@@
$\mathcal{Q}^{\alpha}_{N}(S_{N})\coloneqq\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}\}$
is positivity preserving. Indeed, this follows from the facts that the intersection of two positivity preserving spaces is positivity preserving, and
$\mathcal{Q}^{\alpha}_{N}(\Gamma;S_{N})=\mathcal{Q}^{\alpha}_{N}(S_{N})\cap\mathcal{Q}_{N}(\Gamma;S_{N}).$
To prove that $\mathcal{Q}^{\alpha}_{N}(S_{N})$ is positivity preserving, we first claim that
$\mathcal{Q}^{\alpha}_{N}(S_{N})=\{\Psi\in\mathrm{H}^{1}(S_{N}):\Psi\rvert_{\Gamma_{\rm int}}=0\mbox{ and }\Psi(0,x^{\prime})-\alpha(-1)^{N-1}\Psi(x^{\prime},1)=0\quad\mbox{for a.e. $x^{\prime}\in S_{N-1}$}\}.$ (4.13)
@@
and therefore $\mathcal{Q}^{\alpha}_{N}(S_{N})$ is positivity preserving.
To prove the claim, we first note that, by the definition of $\mathrm{H}^{1}_{\alpha}(I_{N})$ and antisymmetry, any $\Psi\in\mathcal{Q}_{N}^{\alpha}$ satisfies
@@
for any $x^{\prime}=(x_{1},...,x_{N-1})\in I_{N-1}$. This proves the inclusion $\subset$ in (4.13). To prove the opposite inclusion, we observe that for any $\Psi\in\mathrm{H}^{1}(S_{N})$ satisfying $\Psi\rvert_{\Gamma_{\rm int}}=0$ and $\Psi(0,x^{\prime})=(-1)^{N-1}\alpha\Psi(x^{\prime},1)$ for a.e. $x^{\prime}\in S_{N-1}$,
$(T\Psi)(0,\sigma x^{\prime})$ $=\operatorname{sgn}(\sigma)\Psi(0,x^{\prime})=\operatorname{sgn}(\sigma)(-1)^{N-1}\alpha\Psi(x^{\prime},1)=(-1)^{N-1}\alpha(T\Psi)(\sigma x^{\prime},1)$
@@
for any $\sigma\in\mathcal{P}_{N-1}$ and a.e. $x^{\prime}\in S_{N-1}$. As $\{\sigma(S_{N-1})\}_{\sigma\in\mathcal{P}_{N-1}}$ covers $I_{N-1}$ up to finitely many hyperplanes of dimension N2N-2, equation (4.14) holds for almost every $x^{\prime}\in I_{N-1}$. Thus $T\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N})$ and satisfies $(T\Psi)\rvert_{S_{N}}=\frac{1}{\sqrt{N!}}\Psi$, which proves the inclusion $\supset$ in (4.13) and completes the proof. ∎
@@
Let $\alpha\in\mathbb{R}\setminus\{0\}$ and $v\in\mathcal{V}$, then by Lemma A.5, the eigenfunctions of the operator $H_{N}^{\alpha}(v,0)$ (defined in Theorem 2.3) are given by linear combination of Slater determinants of the eigenfunctions of the single-particle operator $h_{\alpha}(v)\coloneqq H_{1}^{\alpha}(v,0)$. In particular, if we denote by $\{\psi_{j}\}_{j\in\mathbb{N}}$ and $\{\lambda_{j}\}_{j\in\mathbb{N}}$ the eigenfunctions and eigenvalues of the operator $h_{\alpha}(v)$ ordered in non-decreasing order, then $H_{N}(v,0)$ has a non-degenerate ground-state if and only if $\lambda_{N}<\lambda_{N+1}$. Thus by Theorem 2.3, we have $\lambda_{N}<\lambda_{N+1}$ for any $N\in\mathbb{N}$ satisfying (2.5), in particular, for N=2k1N=2k-1 if $\alpha\geq 0$ and $N=2k$ for $\alpha\leq 0$. This proves (2.8). The cases of local (or separable) boundary conditions in (2.9) follows from the same argument.